A beaker with 2.00×102 mL of an acetic acid buffer with a pH of 5.000 is sitting on a benchtop. The total molarity of acid and conjugate base in this buffer is 0.100 M. A student adds 8.00 mL of a 0.310 M HCl solution to the beaker. How much will the pH change? The pKa of acetic acid is 4.740.
Answer #1

millimoles of acid + conjugate base = 2.00 x 10^2 x 0.1 =20 ------------->1

pH = pKa + log [conjugate base / acid ]

5.00 = 4.74 + log [conjugate base / acid ]

1.82 = conjugate base / acid

conjugate base = 1.82 acid -------------> 2

from 1 and 2

1.82 acid + acid = 20

2.82 acid = 19

acid = 7.09

conjugate base = 12.9

millimoles of strong acid = C = 8 x 0.310 = 2.48

on additon of C millimoles of acid to buffer

pH = pKa + log [conjugate base - C / acid + C]

pH = 4.74 + log (12.9 - 2.48 / 7.09+ 2.48)

pH = 4.777

pH chnage = 4.777 - 5.000

= -0.223

pH chnage  = - 0.223